### v1

@@ 3.4. $\mathrm{L}^{p}$-based Sobolev regularity of the Bochner Laplacian
Since $\mathrm{W}^{2,p^{*}}_{\#}(\Gamma)\subseteq\mathrm{W}^{1,p^{*}}_{\#}(\Gamma)$, integration-by-parts formula [10, Lemma 1] (valid by density) lets us write
$\int_{\Gamma}u\,\Delta_{\Gamma}v=0,\qquad\forall v\in\mathrm{W}^{2,p^{*}}_{\#}(\Gamma).$
or equivalently, since $\Delta_{\Gamma}$ is an isomorphism from $\mathrm{W}^{2,p^{*}}_{\#}(\Gamma)$ onto $\mathrm{L}^{p^{*}}_{\#}(\Gamma)$ (cf. Lemmas 3.1, 3.2 and 3.3),
$\int_{\Gamma}u\,w=0,\qquad\forall w\in\mathrm{L}^{p^{*}}_{\#}(\Gamma).$
Hence, $u=0$ a.e. on $\Gamma$. ∎
Before continuing, we mention that from a straightforward adaptation of [10, Lemma 2] using 2.10 (compact Sobolev embedding) we deduce that for each $1<p<\infty$, the map $v\mapsto\|\nabla_{\Gamma}v\|_{0,p;\Gamma}$ is an equivalent norm in $\mathrm{W}^{1,p}_{\#}(\Gamma)$. For the proof of the inf-sup condition for (3.42), we adapt the contradiction argument employed in [70].
Proposition 3.13 (inf-sup condition for (3.42)).
Let $1<p<2$ and let $p^{*}:=\frac{p}{p-1}\in(2,\infty)$ be the dual exponent of $p$. Assume $\Gamma$ is of class $C^{3}$. Then, there exists a positive constant $\alpha>0$ such that
$\sup_{0\neq v\in\mathrm{W}^{1,p^{*}}_{\#}(\Gamma)}\frac{\left|\int_{\Gamma}\nabla_{\Gamma}u\cdot\nabla_{\Gamma}v\right|}{\|\nabla_{\Gamma}v\|_{0,p^{*};\Gamma}}\geq\alpha\|\nabla_{\Gamma}u\|_{0,p;\Gamma},\qquad\forall u\in\mathrm{W}^{1,p}_{\#}(\Gamma).$ (3.43)
@@
Step 1: We first establish the (in principle) weaker estimate
$\|u\|_{0,p;\Gamma}+\sup_{0\neq v\in\mathrm{W}^{1,p^{*}}_{\#}(\Gamma)}\frac{\left|\int_{\Gamma}\nabla_{\Gamma}u\cdot\nabla_{\Gamma}v\right|}{\|\nabla_{\Gamma}v\|_{0,p^{*};\Gamma}}\geq\alpha\|\nabla_{\Gamma}u\|_{0,p;\Gamma},\qquad\forall u\in\mathrm{W}^{1,p}_{\#}(\Gamma),$ (3.44)
which will be later used in Step 2.
Since the Bochner Laplacian operator $\mathbf{P}\operatorname{div}_{\Gamma}\nabla_{\Gamma}$ is a continuous bijection from $\mathbf{W}^{2,p^{*}}_{t}(\Gamma)$ onto $\mathbf{L}^{p^{*}}_{t}(\Gamma)$ (cf. 3.11), we can write
$\|\nabla_{\Gamma}u\|_{0,p;\Gamma}=\sup_{\begin{subarray}{c}\mathbf{w}\in\mathbf{L}^{p^{*}}(\Gamma)\\ \mathbf{w}\neq\mathbf{0}\end{subarray}}\frac{\left|\int_{\Gamma}\nabla_{\Gamma}u\cdot\mathbf{w}\right|}{\|\mathbf{w}\|_{0,p^{*};\Gamma}}\lesssim\sup_{\begin{subarray}{c}\mathbf{w}\in\mathbf{L}^{p^{*}}_{t}(\Gamma)\\ \mathbf{w}\neq\mathbf{0}\end{subarray}}\frac{\left|\int_{\Gamma}\nabla_{\Gamma}u\cdot\mathbf{w}\right|}{\|\mathbf{w}\|_{0,p^{*};\Gamma}}\lesssim\sup_{\begin{subarray}{c}\mathbf{v}\in\mathbf{W}^{2,p^{*}}_{t}(\Gamma)\\ \mathbf{v}\neq\mathbf{0}\end{subarray}}\frac{\left|\int_{\Gamma}\nabla_{\Gamma}u\cdot\mathbf{P}\operatorname{div}_{\Gamma}\nabla_{\Gamma}\mathbf{v}\right|}{\|\mathbf{v}\|_{2,p^{*};\Gamma}},\quad\forall u\in\mathrm{W}^{1,p}_{\#}(\Gamma).$ (3.45)
From 2.1 with $\phi\leftarrow u$ and $\mathbf{v}\leftarrow\mathbf{v}$ (valid by density) and integrating by parts, we easily see that
@@
$\left|\int_{\Gamma}\nabla_{\Gamma}u\cdot\mathbf{P}\operatorname{div}_{\Gamma}\nabla_{\Gamma}\mathbf{v}\right|\leq\left|\int_{\Gamma}\nabla_{\Gamma}u\cdot\nabla_{\Gamma}\operatorname{div}_{\Gamma}\mathbf{v}\right|+C\|u\|_{0,p;\Gamma}\|\mathbf{v}\|_{1,p^{*};\Gamma}\qquad\forall u\in\mathrm{W}^{1,p}(\Gamma),\mathbf{v}\in\mathbf{W}^{2,p^{*}}_{t}(\Gamma).$
Plugging this estimate in the right-hand side of (3.45) yields
$\|\nabla_{\Gamma}u\|_{0,p;\Gamma}\lesssim\sup_{\begin{subarray}{c}\mathbf{v}\in\mathbf{W}^{2,p^{*}}_{t}(\Gamma)\\ \operatorname{div}_{\Gamma}\mathbf{v}\neq 0\end{subarray}}\frac{\left|\int_{\Gamma}\nabla_{\Gamma}u\cdot\nabla_{\Gamma}\operatorname{div}_{\Gamma}\mathbf{v}\right|}{\|\mathbf{v}\|_{2,p^{*};\Gamma}}+\|u\|_{0,p;\Gamma}\lesssim\sup_{\begin{subarray}{c}\mathbf{v}\in\mathbf{W}^{2,p^{*}}_{t}(\Gamma)\\ \operatorname{div}_{\Gamma}\mathbf{v}\neq 0\end{subarray}}\frac{\left|\int_{\Gamma}\nabla_{\Gamma}u\cdot\nabla_{\Gamma}\operatorname{div}_{\Gamma}\mathbf{v}\right|}{\|\operatorname{div}_{\Gamma}\mathbf{v}\|_{1,p^{*};\Gamma}}+\|u\|_{0,p;\Gamma}\\ \lesssim\sup_{\begin{subarray}{c}v\in\mathrm{W}^{1,p^{*}}_{\#}(\Gamma)\\ v\neq 0\end{subarray}}\frac{\left|\int_{\Gamma}\nabla_{\Gamma}u\cdot\nabla_{\Gamma}v\right|}{\|v\|_{1,p^{*};\Gamma}}+\|u\|_{0,p;\Gamma},\qquad\forall u\in\mathrm{W}^{1,p}_{\#}(\Gamma),$
which is precisely (3.44).
Step 2: With (3.44) at hand, we now proceed to prove (3.43) by contradiction in the spirit of Peetre–Tartar’s Lemma [27, Lemma A.38]. In fact, assume there exists a sequence $\{u_{n}\}_{n\in\mathrm{N}}\subseteq\mathrm{W}^{1,p}_{\#}(\Gamma)$ such that $\|\nabla_{\Gamma}u_{n}\|_{1,p;\Gamma}=1$ for each $n\in\mathrm{N}$ and
$\sup_{0\neq v\in\mathrm{W}^{1,p^{*}}_{\#}(\Gamma)}\frac{\left|\int_{\Gamma}\nabla_{\Gamma}u_{n}\cdot\nabla_{\Gamma}v\right|}{\|\nabla_{\Gamma}v\|_{0,p^{*};\Gamma}}\xrightarrow[]{n\to\infty}0.$
In particular,
$\lim_{n\to\infty}\int_{\Gamma}\nabla_{\Gamma}u_{n}\cdot\nabla_{\Gamma}v=0,\qquad\forall v\in\mathrm{W}^{1,p^{*}}_{\#}(\Gamma).$ (3.46)
Since $\{u_{n}\}_{n\in\mathrm{N}}$ is bounded in $\mathrm{W}^{1,p}_{\#}(\Gamma)$ (which is reflexive), the Banach–Alaoglu Theorem [12, Theorem 3.18] and 2.10 (compact Sobolev embedding) implies the existence of $u\in\mathrm{W}^{1,p}_{\#}(\Gamma)$ such that (up to a subsequence)
$u_{n}$ $\rightharpoonup u,\qquad\text{in $\mathrm{W}^{1,p}_{\#}(\Gamma)$},$
$u_{n}$ $\to u,\qquad\text{in $\mathrm{L}^{p}(\Gamma)$}.$
Thus, for each $v\in\mathrm{W}^{1,p^{*}}_{\#}(\Gamma)$,
[…]
By 3.12 (injectivity of $\Delta_{\Gamma}$ in $\mathrm{W}^{1,p}_{\#}(\Gamma)$), we deduce that $u=0$ a.e. on $\Gamma$. Hence, passing to the limit $n\to\infty$ in inequality (3.44) for $u=u_{n}$ yields $\lim_{n\to\infty}\|\nabla_{\Gamma}u_{n}\|_{0,p;\Gamma}=0$, which contradicts the fact that $\|\nabla_{\Gamma}u_{n}\|_{0,p;\Gamma}=1$ for each $n\in\mathrm{N}$. This concludes the proof. ∎
Theorem 3.14 (well-posedness of the Laplace–Beltrami operator in $\mathrm{W}^{1,p}_{\#}(\Gamma)$ ($1<p<2$)).
Let $1<p<2$ and let $p^{*}:=\frac{p}{p-1}\in(2,\infty)$ be the dual exponent of $p$. Assume $\Gamma$ is of class $C^{3}$. Then for each $f\in(\mathrm{W}^{1,p^{*}}_{\#}(\Gamma))^{\prime}$, there is a unique $u\in\mathrm{W}^{1,p}_{\#}(\Gamma)$ solution of (3.42). Moreover, there exists a positive constant $C$, depending only on $\Gamma$ and $p$ such that

### v2

@@ 4. Tangent Stokes equations
Fix $\uppi\in\mathrm{L}_{\#}^{p}(\Gamma)$. By the higher regularity result within 2.15 (Laplace–Beltrami) we know there exists a unique $\phi\in\mathrm{W}^{2,p^{*}}_{\#}(\Gamma)$ such that $\Delta_{\Gamma}\phi=|\uppi|^{p-2}\uppi-\alpha_{\uppi}$ a.e. on $\Gamma$, where $\alpha_{\pi}:=\frac{1}{|\Gamma|}\int_{\Gamma}|\uppi|^{p-2}\uppi$. Furthermore, $\|\phi\|_{\mathrm{W}^{2,p^{*}}(\Gamma)}\lesssim\||\uppi|^{p-2}\uppi-\alpha_{\uppi}\|_{0,p^{*};\Gamma}\lesssim\|\uppi\|_{0,p;\Gamma}^{p-1}$. We use the function $\mathbf{w}:=\nabla_{\Gamma}\phi\in\mathbf{W}^{1,p^{*}}_{t}(\Gamma)$, that satisfies $\operatorname{div}_{\Gamma}\mathbf{w}=|\uppi|^{p-2}\uppi-\alpha_{\uppi}$ on $\Gamma$ and $\|\mathbf{w}\|_{1,p^{*};\Gamma}\lesssim\|\phi\|_{\mathrm{W}^{2,p^{*}}(\Gamma)}\lesssim\|\uppi\|_{0,p;\Gamma}^{p-1}$, to bound the supremum from below. This yields
$\sup_{\begin{subarray}{c}\mathbf{v}\in\mathbf{W}^{1,p^{*}}_{t}(\Gamma)\\ \mathbf{v}\neq\mathbf{0}\end{subarray}}\frac{\left|\int_{\Gamma}\uppi\operatorname{div}_{\Gamma}\mathbf{v}\right|}{\|\mathbf{v}\|_{1,p^{*};\Gamma}}\geq\frac{\int_{\Gamma}\uppi\operatorname{div}_{\Gamma}\mathbf{w}}{\|\mathbf{w}\|_{1,p^{*};\Gamma}}\gtrsim\frac{\|\uppi\|_{0,p;\Gamma}^{p}-\alpha_{\uppi}\int_{\Gamma}\pi}{\|\uppi\|_{0,p;\Gamma}^{p-1}}=\|\uppi\|_{0,p;\Gamma},$
which finishes the proof. ∎
Recalling 3.3 (Poincaré inequality), we are now in position of establishing the well-posedness of formulation (1.4) (equivalently, (4.1) for $p=2$), which follows from the classical Babuška–Brezzi theory.
Theorem 4.2 (well-posedness of tangent Stokes in $\mathbf{H}^{1}_{t}(\Gamma)\times\mathrm{L}^{2}_{\#}(\Gamma)$).
Assume $\Gamma$ is of class $C^{2}$. For each $\mathbf{f}\in(\mathbf{H}^{1}_{t}(\Gamma))^{\prime}$ and $g\in\mathrm{L}_{\#}^{2}(\Gamma)$, there exists a unique $(\mathbf{u},\uppi)\in\mathbf{H}^{1}_{t}(\Gamma)\times\mathrm{L}_{\#}^{2}(\Gamma)$ solution of (4.1) for $p=2$. Moreover, there is a constant $C>0$, depending only on $\Gamma$, such that
$\|\mathbf{u}\|_{1,\Gamma}+\|\uppi\|_{0,\Gamma}\leq C\left(\|\mathbf{f}\|_{(\mathbf{H}^{1}_{t}(\Gamma))^{\prime}}+\|g\|_{0,\Gamma}\right).$
@@
Recalling the Poincaré inequality (3.7) for $p=2$, the bilinear form $(\mathbf{v},\mathbf{w})\mapsto\int_{\Gamma}\nabla_{\Gamma}\mathbf{v}:\nabla_{\Gamma}\mathbf{w}$ is coercive in $\mathbf{H}^{1}_{t}(\Gamma)$. Hence, in view of the inf-sup condition (4.2) for $p=2$, the assertion follows from a straightforward application of the Babuška–Brezzi theory in Hilbert spaces [29, Theorem 2.1] to problem (1.4). ∎
The well-posedness of (4.1) for general $p\in(1,\infty)$ is reserved to Section 4.2, for which the tools developed in the forthcoming Section 4.1 are essential.
4.1. Helmholtz–Weyl decomposition of $\mathbf{W}^{m,p}_{t}(\Gamma)$
Assume $\Gamma$ is of class $C^{m,1}$ for some nonnegative integer $m$. We define
@@ 4.1. Helmholtz–Weyl decomposition of $\mathbf{W}^{m,p}_{t}(\Gamma)$
$\langle\operatorname{div}_{\Gamma}\mathbf{v},q\rangle_{(\mathrm{W}^{1,p^{*}}(\Gamma))^{\prime}\times\mathrm{W}^{1,p^{*}}(\Gamma)}=-\int_{\Gamma}\mathbf{v}\cdot\nabla_{\Gamma}q,\qquad\forall q\in\mathrm{W}^{1,p^{*}}(\Gamma).$ (4.3)
Notice that for $m\geq 1$, $\operatorname{div}\mathbf{v}$ as an element of $(\mathrm{W}^{1,p^{*}}(\Gamma))^{\prime}$ is identified with the function $\operatorname{div}\mathbf{v}\in\mathrm{W}^{m-1,p}_{\#}(\Gamma)$ via
$\langle\operatorname{div}_{\Gamma}\mathbf{v},q\rangle_{(\mathrm{W}^{1,p^{*}}(\Gamma))^{\prime}\times\mathrm{W}^{1,p^{*}}(\Gamma)}=\int_{\Gamma}q\,\operatorname{div}_{\Gamma}\mathbf{v},\qquad\forall q\in\mathrm{W}^{1,p^{*}}(\Gamma),$
because of integration-by-parts formula (2.5) and a density argument. Consequently, the condition $\operatorname{div}_{\Gamma}\mathbf{v}=0$ in the definition of $\mathbf{W}^{m,p}_{t,\sigma}(\Gamma)$ for $m\geq 1$ can be understood to be pointwise a.e.
Proposition 4.3 (Helmholtz–Weyl decomposition on $\Gamma$).
Let $p\in(1,\infty)$. Let $m$ be a nonnegative integer and assume $\Gamma$ is of class $C^{1}$ if $m=0$ or $C^{m,1}$ if $m\geq 1$. Then $\mathbf{W}^{m,p}_{t}(\Gamma)$ can be decomposed as
Furthermore, this decomposition is stable in the sense that there exists a positive constant $C>0$, depending only on $m$, $p$ and $\Gamma$, such that for each $\mathbf{v}=\mathbf{v}_{1}+\mathbf{v}_{2}\in\mathbf{W}^{m,p}_{t}(\Gamma)$ with $\mathbf{v}_{1}\in\mathbf{W}^{m,p}_{t,\sigma}(\Gamma)$ and $\mathbf{v}_{2}\in\nabla_{\Gamma}\mathrm{W}^{m+1,p}_{\#}(\Gamma)$ it holds that
$\|\mathbf{v}_{1}\|_{m,p;\Gamma}+\|\mathbf{v}_{2}\|_{m,p;\Gamma}\leq C\|\mathbf{v}\|_{m,p;\Gamma}.$
For $p=2$, this decomposition is $\mathrm{L}^{2}$-orthogonal.
Finally, if $m=0$ and $\Gamma$ is only of class $C^{0,1}$, then the stable decomposition (4.4) still holds as long as $p\in(2-\varepsilon,2+\varepsilon)$, for some $\varepsilon>0$ depending only on $\Gamma$.